# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v1

@@ V-C Can Our Method Reach Goals?
TABLE III: Breakdown of the average runtime (wall clock) for Algorithm 2. IGH stands for Initial Guess Heuristic. Each cell displays the mean (standard deviation).
VI Stabilization under Second-Order Dynamics
In Section V, we presented results for running MPC on the CQDC dynamics (8). However, although the CQDC dynamics removes the force-at-a-distance effect from smoothing, it is still different from real contact physics: not only are second-order effects ignored, it also introduces a small gap between objects undergoing sliding friction [10, Section IV-A2], which is an artifact known as “hydroplaning” [67, Appendix B] shared by contact dynamics formulations that utilize Anitescu’s convex relaxation for contact dynamics [51].
@@ VI Stabilization under Second-Order Dynamics
VI-A Implementation Heuristics for Second-order Dynamics
@@
Firstly, instead of running MPC as quickly as possible, we compute a longer trajectory (usually 1-2 seconds) using Algorithm 2 for multiple steps ($H>1$) (Algorithm 3), roll out $L_{u}$ in its entirety on the second-order dynamics (Algorithm 3), and repeat. Although we can re-plan at a higher rate by simply replacing the CQDC dynamics ($f$ in Algorithm 2 of Algorithm 2) with $f_{\text{2nd}}$, we observed that doing high-rate MPC without waiting for the system to stabilize is inconsistent with the quasi-dynamic nature of the CQDC model. In practice, we found CQDC-based MPC works better when we let the system stabilize to an equilibrium before re-planning.
Secondly, we apply the initial guess heuristics more frequently to compensate for the side effects of the hydroplaning artifact. When running trajectories generated by Algorithm 2 open-loop on second-order dynamics, we observed that the robot sometimes loses contact with the object. We hypothesized that this is caused by the hydroplaning artifact of the CQDC dynamics, which allows the robot to exert sliding friction forces on the object without actually touching it. As the hydroplaning artifact does not exist for sticking friction [51], we can alleviate hydroplaning by regularly pulling the robot back into contact with object using the initial guess heuristics (Section IV-B). Specifically, in Algorithm 2, instead of only applying the initial guess heuristics at the first iteration (Algorithm 2 to Algorithm 2), we apply it at every iteration to the corresponding $q_{t}^{\mathrm{a}}$.
@@
1 Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, planning horizon $T$, iterations limit $n_{\mathrm{max}}$, MPC rollout horizon $H$, number of re-plans $N$;
@@ 7       $i\leftarrow i+1$
(a) Iiwa Bimanual
(b) Allegro Hand
Figure 11: Results for running the three variants of Algorithm 3, namely open-loop, no heuristics and close-loop, on the pairs of goals and initial conditions generated in Section V-A1. All figure elements follow the same definitions used in Figure 9.
VI-B Experiment Setup
We evaluate Algorithm 3 using the same final-distance-to-goal metric proposed in Section V-A2 on both IiwaBimanual and AllegroHand. Hyperparameter choices are summarized in Section X-E. For each system, we run two ablation variants of our algorithm, each omitting a crucial component.
  • Closed-loop. This is running Algorithm 3 as is with multiple re-plans ($N>1$).
  • No Heuristics. We replace $\mathbf{MPC}_{\text{proj}}$ with the $\mathbf{MPC}$ defined in Algorithm 2. In other words, we apply the initial guess heuristics only at the first iteration in MPC, instead of at every iteration.
  • Open-loop. This can be interpreted as reducing feedback rate to the extreme: there is no re-planning at all. Instead, we plan a long trajectory that gets the object all the way to the goal and run it open-loop.
Simulation Setup. We incorporate the IiwaBimanual and AllegroHand systems presented in Section V into Drake [15], which employs a complete second-order dynamics model along with state-of-the-art contact solvers [68] that do not have the hydroplaning effect. We maintain identical collision geometries, robot controller stiffness, and friction coefficients between the CQDC dynamics and Drake.
Hardware Setup. We keep the MPC and physical parameters of the system, such as robot feedback gains, object shape, mass, and friction, consistent between Drake and hardware.
We measure the object pose utilizing the OptiTrack motion capture system. Passive spherical reflectors are sufficient for the bucket in IiwaBimanual. For the cube in AllegroHand, in order to not alter the collision geometry with external markers, we embed light-emitting active markers in the cube. In addition, we post-process the cube pose by solving an inverse kinematics problem that projects the raw OptiTrack measurement out of penetration with the Allegro hand.
VI-C Results & Discussion
We plot the results of our experiments in Figure 11, and display the translation and rotation errors separately in Table IV. We discuss some of our findings from the experiments.
@@ VI-C Results & Discussion
(0.065) (0.079) (0.009) (0.197)
@@
TABLE IV: Translation and rotation errors for simulation (Sim) and hardware (HW) experiments. Each cell displays the mean (standard deviation).
VI-C1 Closed-loop vs. Open-loop
@@
Challenges due to slipping is also evident in the distribution of tracking errors in the scatter plots in Figure 11b. Under the closed-loop MPC, Allegro’s tracking error remains low for large yaw angles when pitch angle is near zero, because the cube’s bottom face stays close the palm for such goals. In other goal orientations requiring more lift, the cube is less supported by the palm, increasing the risk of losing grip.
VI-C3 Closed-loop vs. No Heuristics
@@
In fact, for each pair of iniditial and goal configurations, we can compute the path length travelved by the robot as follows:
Robot Path $\text{Robot Path Length}\;=\;\int_{0}^{t_{\text{final}}}\|\dot{q}^{\mathrm{a}}(t)\|_{2}\,\mathrm{d}t,$ (39)
@@
VI-C4 Hardware Experiments
Using voxel-grid down-sampling, we get 50 uniformly-spaced pairs of initial states and goals for each system from the pairs used for simulation experiments. Shown in the last two rows of Table IV, the average errors of hardware experiments are comparable to those of simulation. All hardware and simulation videos are available in an interactive plot on our project website.
@@ VI-C4 Hardware Experiments
VII Generating Goal-conditioned Contact Configurations
Through our method presented in Section IV, as well as the results in Section V and Section VI, we have shown how our proposed MPC allows successful stabilization to local goals.
@@ VII-A Problem Specification
VII-A1 Finite-Horizon Value Function of the MPC Policy
@@
We note that although MPC rollout in Algorithm 2 accepts a full configuration $q_{g}$ as a goal, we can give it unactuated configurations only ($q^{\mathrm{o}}_{g}$) by setting the cost terms of the actuated objects to zero, $\mathbf{Q}_{a}=\mathbf{0}$, as we did in Section V-A2. In addition, due to the efficiency of the MPC controller, the finite-hozion value function can be quickly queried online. However, the landscape of this value function, as visualized for a simple problem in Figure 16, is multi-modal with many local minima and maxima. This hints at the necessity of global optimization when we search for the minimizers of (41).
Figure 16: Visualization of the landscape of (41) for the planar pusher-T example inspired by [69]. The black T corresponds to the current object configuration ${q^{\mathrm{o}}}$, and the green T the goal object configuration $q^{\mathrm{o}}_{g}$. The landscape corresponds to the position of the round pusher $q^{\mathrm{a}}$, colored with the cost function $V({q^{\mathrm{a}}};{q^{\mathrm{o}}},q^{\mathrm{o}}_{g})$. The red pusher configuration corresponds to the global optima of the landscape.
VII-A2 Robustness Regularizer
@@
For example, consider the task of pushing the ball slightly to the right in Figure 17. Both configurations are nearly equal in the goal-reaching cost (see Table V); yet, the configuration in Figure 17a would be preferred in practice, as both fingers can be used to reject small disturbances.
Figure 17: Top Row: Comparison of two configurations (a) and (b) for the 2D planar hand system in Figure 5. Although both configurations are advantageous for moving the object to the goal, they share the same MPC value function cost (Table V). Bottom Row: Plot of the wrench set (illustrated in red) and the maximum inscribed sphere (blue), where (c) corresponds to configuration (a) and (d) corresponds to (b). Note that configuration (a) has a much larger inscribed sphere due to the anti-podal grasp.
Configuration (a) (b)
MPC Value function $V$ 2.82e-4 2.16e-4
Max inscribed sphere radius. 7.12e-3 1.68e-3
TABLE V: Comparison of MPC value function cost and maximum inscribed sphere radius for two configurations in Figure 17.
To formalize this notion of robustness, we take inspiration from classical grasping metrics [61, 70, 71], as well as the connection between the RA-CTR and the classical wrench set in Section III-E. Specifically, we adopt a worst-case metric [61] that reasons about the maximum wrench a grasp can resist along any direction. Geometrically, this quantity corresponds to the maximum-inscribed sphere in the wrench set.
@@
VII-A3 Total Cost
We combine the two costs in Section VII-A1 and Section VII-A2 use a weighting term $\alpha\in\mathbb{R}_{\geq 0}$:
@@ VII-A3 Total Cost
VII-B Solving (40) by Sampling-based Optimization
Solving (40) is challenging, as i) the gradient of the cost function is quite difficult to obtain [71], and ii) it requires global search, as evidenced by the nonconvex cost landscape in Figure 16. As a result, we resort to a simple sampling-based search which samples from the feasible set of (40) by rejection sampling, then chooses the best sample.
However, due to the high variance of this process in high-dimensional spaces, as well as the complexity of navigating high-dimensional configuration spaces, we found that directly applying this strategy is not very effective beyond simple planar problems. To scale the sampling-based search to AllegroHand, we introduce a few heuristic changes to make this optimization more tractable, which we detail in Section X-F.
VII-C Results
Results in Table VI indicate that our method achieves error on the order of $10$mm in position, and $30$ mrad ($1.7^{\circ}$) in orientation. Moreover, the discovered initial hand configurations in Figure 18 agree with our intuition on how the cube should be grasped if we want to move the cube towards the goals.
We note that the given result was obtained with $1000$ samples on the reduced-order model. Reducing the number of samples for lower computation time is possible, but would result in higher variance in performance. In addition, we found that other types of motions, such as pitch $0^{\circ}$ to $-90^{\circ}$, or rotation along the $x$-axis (roll), are physically very difficult on the Allegro hand.
@@ VII-C Results
Yaw $0^{\circ}$ to $45^{\circ}$ $12.05$ $(3.49)$ $32.41$ $(27.14)$ $57.83$ $(2.08)$
Yaw $45^{\circ}$ to $90^{\circ}$ $9.85$ $(10.88)$ $15.72$ $(21.08)$ $57.28$ $(0.72)$
Pitch $0^{\circ}$ to $90^{\circ}$ $21.42$ $(8.91)$ $32.12$ $(32.03)$ $57.84$ $(0.73)$
TABLE VI: Performance of our algorithm for finding initial actuator configurations. Results are obtained with $10$ repeated runs. Each run is optimized using $1000$ samples. Each cell displays the mean (standard deviation).
VIII Global Planning with Roadmaps
Using the method for generating contact configurations in Section VII, we present a simple recipe for global search, in which we chain local plans together to efficiently reach global goals which are challenging for local MPC. Our method, inspired by the Probabilistic Roadmap (PRM)[72], consists of an offline phase in which the roadmap is constructed, and an online phase where we reach arbitrary goals using the roadmap.
VIII-A Roadmap Construction

### v3

@@ 5.3.3 Computation Time
Table 5: Breakdown of the average runtime (wall clock) for Algorithm 2. IGH stands for Initial Guess Heuristic. Each cell displays the mean (standard deviation).
6 Stabilization under Second-Order Dynamics
In Section 5, we presented results for running MPC on the CQDC dynamics (8). However, although the CQDC dynamics removes the force-at-a-distance effect from smoothing, it is still different from real contact physics: not only are second-order effects ignored, it also introduces a small gap between objects undergoing sliding friction (Pang and Tedrake, 2021, Section IV-A2), which is an artifact known as “hydroplaning” (Tedrake, 2023, Appendix B) shared by contact dynamics formulations that utilize Anitescu’s convex relaxation for contact dynamics (Anitescu, 2006).
@@ 6 Stabilization under Second-Order Dynamics
6.1 Implementation Heuristics for Second-order Dynamics
@@
Firstly, instead of running MPC as quickly as possible, we compute a longer trajectory (usually 1-2 seconds) using Algorithm 2 for multiple steps ($H>1$) (Algorithm 3), roll out $L_{u}$ in its entirety on the second-order dynamics (Algorithm 3), and repeat. Although we can re-plan at a higher rate by simply replacing the CQDC dynamics ($f$ in Algorithm 2 of Algorithm 2) with $f_{\text{2nd}}$, we observed that doing high-rate MPC without waiting for the system to stabilize is inconsistent with the quasi-dynamic nature of the CQDC model. In practice, we found CQDC-based MPC works better when we let the system stabilize to an equilibrium before re-planning in order to ensure $f$ remains similar to $f_{\text{2nd}}$. See Section A.6 for more details about the effect of $H$ on closed-loop tracking performance.
Secondly, we apply the initial guess heuristics more frequently to compensate for the side effects of the hydroplaning artifact. When running trajectories generated by Algorithm 2 open-loop on second-order dynamics, we observed that the robot sometimes loses contact with the object. We hypothesized that this is caused by the hydroplaning artifact of the CQDC dynamics, which allows the robot to exert sliding friction forces on the object without actually touching it. As the hydroplaning artifact does not exist for sticking friction (Anitescu, 2006), we can alleviate hydroplaning by regularly pulling the robot back into contact with object using the initial guess heuristics (Section 4.2). Specifically, in Algorithm 2, instead of only applying the initial guess heuristics at the first iteration (Algorithm 2 to Algorithm 2), we apply it at every iteration to the corresponding $q_{t}^{\mathrm{a}}$.
@@
1 Input: Initial state $q_{0}$, goal state $q_{\text{goal}}$, planning horizon $T$, iterations limit $n_{\mathrm{max}}$, MPC rollout horizon $H$, number of re-plans NN;
@@ 7 $i\leftarrow i+1$
(a) IiwaBimanual: implementation heuristics ablation, all using R-CTR.
(b) IiwaBimanual: closed-loop trust region comparison. R-CTR results are the same as the closed-loop results in (a).
(c) AllegroHand: implementation heuristics ablation, all using R-CTR.
(d) AllegroHand: closed-loop trust region comparison. R-CTR results are the same as the closed-loop results in (c).
Figure 12: Results for running the five variants of Algorithm 3, namely open-loop, no heuristics, close-loop, and closed-loop with CTR and ETR, on the pairs of goals and initial conditions generated in Section 5.1.1. All figure elements follow the same definitions used in Figure 11.
6.2 Experiment Setup
We evaluate Algorithm 3 using the same final-distance-to-goal metric proposed in Section 5.1.2 on both IiwaBimanual and AllegroHand. Hyperparameter choices are summarized in Section A.6. For each system, we run four ablation variants of our algorithm, each omitting a crucial component.
  • Closed-loop. This is running Algorithm 3 as is with multiple re-plans ($N>1$) and R-CTR.
  • No Heuristics. We keep using R-CTR, but replace $\mathbf{MPC}_{\text{proj}}$ with the $\mathbf{MPC}$ defined in Algorithm 2. In other words, we apply the initial guess heuristics only at the first iteration in MPC, instead of at every iteration.
  • Open-loop. We keep using R-CTR, but reduce feedback rate to the extreme: there is no re-planning at all. Instead, we plan a long trajectory that gets the object all the way to the goal and run it open-loop.
  • Closed-loop with CTR/ETR. We study how the other trust region variants affect closed-loop performance.
Simulation Setup. We incorporate the IiwaBimanual and AllegroHand systems presented in Section 5 into Drake (Tedrake and the Drake Development Team, 2019), which employs a complete second-order dynamics model along with state-of-the-art contact solvers (Castro et al., 2023) that do not have the hydroplaning effect. We maintain identical collision geometries, robot controller stiffness, and friction coefficients between the CQDC dynamics and Drake.
Hardware Setup. To minimize the sim-to-real gap, we carefully aligned key parameters between our simulation model in Drake and the physical system. Specifically, we ensured a close correspondence for: (i) the collision geometries used for planning, (ii) the feedback gains of the robot’s low-level controllers, and (iii) the inertial and friction coefficients. This alignment was crucial for the successful transfer of our method to hardware.
As the focus of this work is state-based planning and control for contact-rich manipulation, we measure the object pose utilizing the OptiTrack motion capture system. Passive spherical reflectors are sufficient for the bucket in IiwaBimanual. For the cube in AllegroHand, in order to not alter the collision geometry with external markers, we embed infrared light-emitting active markers in the cube In order to minimize the effect of marker occlusion from the hand, we adopt a similar marker placement as the cube used by OpenAI et al. (2019). In addition, we post-process the cube pose by solving an inverse kinematics problem that projects the raw OptiTrack measurement out of penetration with the Allegro hand. In practice, our pose estimation setup proved reliable enough to forgo additional measures for addressing uncertainty.
6.3 Results & Discussion
We plot the results of our experiments in Figure 12, and display the translation and rotation errors separately in Table 6. We discuss some of our findings from the experiments.
@@ 6.3 Results & Discussion
with R-CTR (0.065) (0.079) (0.009) (0.197)
Closed-loop Sim 0.040 0.076 0.013 0.349
with ETR (0.167) (0.214) (0.010) (0.267)
Closed-loop Sim 0.031 0.051 0.013 0.289
with CTR (0.122) (0.164) (0.010) (0.211)
@@
Table 6: Translation and rotation errors for simulation (Sim) and hardware (HW) experiments. Each cell displays the mean (standard deviation).
6.3.1 Closed-loop vs. Open-loop
@@
Challenges due to slipping is also evident in the distribution of tracking errors in the scatter plots in Figure 12c. Under the closed-loop MPC, Allegro’s tracking error remains low for large yaw angles when pitch angle is near zero, because the cube’s bottom face stays close the palm for such goals. In other goal orientations requiring more lift, the cube is less supported by the palm, increasing the risk of losing grip.
6.3.3 Closed-loop vs. No Heuristics
@@
In fact, for each pair of initial and goal configurations, we can compute the path length traveled by the robot as follows:
$\text{Robot Path Length}\;=\;\int_{0}^{t_{\text{final}}}\|\dot{q}^{\mathrm{a}}(t)\|_{2}\,\mathrm{d}t,$ (39)
@@
6.3.4 R-CTR vs. ETR/CTR
The performance comparison between the different trust region formulations under second-order dynamics follows the same general trend observed in our local MPC experiments (Figure 11). R-CTR remains the top performer on both systems, achieving the lowest average error and exhibiting a tighter error distribution with fewer large-error outliers.
On IiwaBimanual, the advantage of R-CTR is more significant, showing a larger improvement over both the more restrictive CTR and the baseline ETR. For AllegroHand, the result is more nuanced: while R-CTR and CTR perform almost identically, both offer a clear improvement over ETR, which produces a visibly wider error distribution.
6.3.5 Hardware Experiments
Using voxel-grid down-sampling, we selected a subset of 50 uniformly-spaced pairs of initial states and goals for each system from the pairs used for simulation experiments. As shown in the final two rows of Table 6, the average errors of hardware experiments are highly comparable to our simulation results. This strong sim-to-real correspondence is expected, given our careful alignment of the simulation and hardware environments. All hardware and simulation videos are available in an interactive plot on our project website.
@@ 6.3.5 Hardware Experiments
7 Generating Goal-conditioned Contact Configurations
Through our method presented in Section 4, as well as the results in Section 5 and Section 6, we have shown how our proposed MPC allows successful stabilization to local goals.
@@ 7.1 Problem Specification
7.1.1 Finite-Horizon Value Function of the MPC Policy
@@
We note that although MPC rollout in Algorithm 2 accepts a full configuration $q_{g}$ as a goal, we can give it unactuated configurations only ($q^{\mathrm{o}}_{g}$) by setting the cost terms of the actuated objects to zero, $\mathbf{Q}_{a}=\mathbf{0}$, as we did in Section 5.1.2. In addition, due to the efficiency of the MPC controller, the finite-hozion value function can be quickly queried online. However, the landscape of this value function, as visualized for a simple problem in Figure 17, is multi-modal with many local minima and maxima. This hints at the necessity of global optimization when we search for minimizers of (41).
Figure 17: Visualization of the landscape of (41) for the planar pusher-T example inspired by Chi et al. (2024). The black T corresponds to the current object configuration $q^{\mathrm{o}}$}, and the green T the goal object configuration $q^{\mathrm{o}}_{g}$. The landscape corresponds to the position of the round pusher $q^{\mathrm{a}}$, colored with the cost function $V({q^{\mathrm{a}}};{q^{\mathrm{o}}},q^{\mathrm{o}}_{g})$. The red pusher configuration corresponds to the global optima of the landscape.
7.1.2 Robustness Regularizer
@@
For example, consider the task of pushing the ball slightly to the right in Figure 18. Both configurations are nearly equal in the goal-reaching cost (see Table 7); yet, the configuration in Figure 18a would be preferred in practice, as both fingers can be used to reject small disturbances.
To formalize this notion of robustness, we take inspiration from classical grasping metrics (Ferrari and Canny, 1992; Dai et al., 2018; Li et al., 2023), as well as the connection between the RA-CTR and the classical wrench set in Section 3.5. Specifically, we adopt a worst-case metric (Ferrari and Canny, 1992) that reasons about the maximum wrench a grasp can resist along any direction. Geometrically, this quantity corresponds to the maximum-inscribed sphere in the wrench set.
@@
7.1.3 Total Cost
We combine the two costs in Section 7.1.1 and Section 7.1.2 use a weighting term $\alpha\in\mathbb{R}_{\geq 0}$:
@@ 7.1.3 Total Cost
Figure 18: Top Row: Comparison of two configurations (a) and (b) for the 2D planar hand system in Figure 5. Although both configurations are advantageous for moving the object to the goal, they share the same MPC value function cost (Table 7). Bottom Row: Plot of the wrench set (illustrated in red) and the maximum inscribed sphere (blue), where (c) corresponds to configuration (a) and (d) corresponds to (b). Note that configuration (a) has a much larger inscribed sphere due to the anti-podal grasp.
Configuration (a) (b)
MPC value function $V$ $2.82\times 10^{-4}$ $2.16\times 10^{-4}$
Max inscribed sphere radius $7.12\times 10^{-3}$ $1.68\times 10^{-3}$
Table 7: Comparison of MPC value function cost and maximum inscribed sphere radius for two configurations in Figure 18.
7.2 Solving (40) by Sampling-based Optimization
Solving (40) is challenging, as i) the gradient of the cost function is quite difficult to obtain (Li et al., 2023), and ii) it requires global search, as evidenced by the nonconvex cost landscape in Figure 17. As a result, we resort to a simple sampling-based search which samples from the feasible set of (40) by rejection sampling, then chooses the best sample.
However, due to the high variance of this process in high-dimensional spaces, as well as the complexity of navigating high-dimensional configuration spaces, we found that directly applying this strategy is not very effective beyond simple planar problems. To scale the sampling-based search to AllegroHand, we introduce a few heuristic changes to make this optimization more tractable, which we detail in Section A.7.
@@ 7.2 Solving (40) by Sampling-based Optimization
Yaw $0^{\circ}\!\to\!45^{\circ}$ 12.05 (3.49) 32.41 (27.14) 57.83 (2.08)
Yaw $45^{\circ}\!\to\!90^{\circ}$ 9.85 (10.88) 15.72 (21.08) 57.28 (0.72)
Pitch $0^{\circ}\!\to\!90^{\circ}$ 21.42 (8.91) 32.12 (32.03) 57.84 (0.73)
Table 8: Performance of our algorithm for finding initial actuator configurations. Results are obtained with $10$ repeated runs. Each run is optimized using $1000$ samples. Each cell displays the mean (standard deviation).
7.3 Results
Results in Table 8 indicate that our method achieves error on the order of 1010mm in position, and $30$ mrad ($1.7^{\circ}$) in orientation. Moreover, the discovered initial hand configurations in Figure 19 agree with our intuition on how the cube should be grasped if we want to move the cube towards the goals.
We note that the given result was obtained with $1000$ samples on the reduced-order model. Reducing the number of samples for lower computation time is possible, but would result in higher variance in performance. In addition, we found that other types of motions, such as pitch $0^{\circ}$ to $90^{\circ}$, or rotation along the $x$-axis (roll), are physically very difficult on the Allegro hand.
8 Global Planning with Roadmaps
Using the method for generating contact configurations in Section 7, we present a simple recipe for global search, in which we chain local plans together to efficiently reach global goals which are challenging for local MPC. Our method, inspired by the Probabilistic Roadmap (PRM)(Choset et al., 2005), consists of an offline phase in which the roadmap is constructed, and an online phase where the same roadmap is reused to reach any new goals.
8.1 Roadmap Construction